### v1

# LauraTSE: Target Speaker Extraction using Auto-Regressive Decoder-Only Language Models

Abstract
We propose LauraTSE, an Auto-Regressive Decoder-Only Language Model for Target Speaker Extraction (TSE) based on the LauraGPT backbone. It employs a small-scale auto-regressive decoder-only language model which takes the continuous representations for both the mixture and the reference speeches and produces the first few layers of the target speech’s discrete codec representations. In addition, a one-step encoder-only language model reconstructs the sum of the predicted codec embeddings using both the mixture and the reference information. Our approach achieves superior or comparable performance to existing generative and discriminative TSE models. To the best of our knowledge, LauraTSE is the first single-task TSE model to leverage an auto-regressive decoder-only language model as the backbone.
Index Terms:
target speaker extraction, auto-regressive decoder-only language models, discrete tokens, neural audio codec
I Introduction
Target Speaker Extraction (TSE) aims at extracting target speaker’s speech from a mixture using auxiliary information like reference speech, spatial information, and visual information etc. regarding the target speaker [1]. Current dominant approaches utilize discriminative models which try to directly map the mixture speech to clean speech [2, 3, 4, 5]. However, this method might struggle for unseen data and sometimes even introduce undesirable distortions [6]. Also, when data is highly corrupted with a low Signal-to-Noise Ration (SNR), directly mapping might not be optimal. Generative models, on the other hand, have gained the attention for its capability in dealing with unseen noises compared with discriminative models [7, 8, 9] as well as its superior performances in terms of the audio quality [10, 11, 12, 13]. Rather than learning the map from noisy speech to clean speech, generative models aim at studying the underlying distribution of the clean output [10]. Generative models like diffusion models [10] and variational autoencoders(VAE) [14] have been studied for TSE. Language models (LMs) are also studied. TSELM utilizes encoder-only language models and discrete tokens from WavLM [15] for TSE [16]. AnyEnhance [12] utilizes a masking encoding language models for multi-task speech processing, including TSE.
However, Auto-Regressive (AR) decoder-only LMs, as another important class of generative models, have not been thoroughly studied for TSE. One of the existing works related is SpeechX [17], which proposes a multi-task speech processing model utilizing an AR decoder-only model. However, this work still has several limitations. Firstly, the TSE task in the paper remains relatively simple, which does not demonstrate the full capability of AR models on TSE tasks. Secondly, this work uses discrete representations as the input to the AR model, which turns out to be suboptimal for certain tasks compared with continuous features as mentioned in [18]. Finally, since SpeechX is designed as a multi-task system—handling tasks such as noise suppression, speech removal, TSE, and TTS—it remains unclear whether a small-scale single-task AR decoder-only model can effectively perform TSE on its own.
To use AR models, we need to discretize audio into tokens for the classification loss. There are two main approaches for audio discretization. The first approach applies Kmeans clustering on the outputs of self-supervised learning (SSL) models, as done in [19, 16, 20, 21]. However, this method has been shown to lose speaker-specific information [16, 21], likely because SSL models are primarily optimized for capturing semantic content rather than speaker characteristics. The second approach leverages neural audio codecs [22, 17, 18], which discretize audio into multiple layers of finite token sequences. This method has shown greater promise in preserving both acoustic and speaker-related information [18], making it a potential for tasks like TSE where speaker information needs to be well-preserved.
Figure 1: Overview of lauraTSE.
In this work, we propose LauraTSE, an AR decoder-only LM for TSE, built upon the LauraGPT backbone [18]. The model takes the log-mel spectrograms of both the reference and mixture speech as input and utilizes a neural audio codec to represent the output of the AR model. LauraTSE consists of two components: (1) an AR decoder-only LM that predicts the first few layers of the codec representations of the target speech, and (2) a one-step encoder-only LM that directly predicts the sum of all layers of the codec embeddings by integrating information from both the mixture and reference signals. An overview of the model architecture is provided in Fig. 1. To the best of our knowledge, we are the first to conduct single-task TSE using AR decoder-only LMs with continuous input features. Experimental results show that LauraTSE achieves performance comparable to or better than existing generative approaches and discriminative approaches. Our demos and code are available at  [^1] https://beilong-tang.github.io/lauraTSE.demo/.
II Method
II-A Encoder
The first stage of LauraTSE is the encoding stage. Similar to the Speech Enhancement (SE) tasks in LauraGPT [18], we begin by computing the log-mel spectrograms for both the reference and mixture speech signals. These spectrograms are then passed through a shared Conformer [23] model, which produces continuous embeddings for the reference and mixture, denoted as $r$ and $m$, respectively. This stage acts as an adapter, transforming both the mixture and reference inputs into a suitable representation space for the subsequent AR decoder-only LM. Notably, unlike SpeechX [17], which utilizes discrete embeddings obtained from a neural audio codec, our approach retains continuous representations learned directly from the task. Our experimental results show that using continuous features as inputs leads to better performance compared to discrete token embeddings.
II-B Auto-Regressive Decoder-Only Language Model
TABLE I: Results on Libri2Mix clean. In the ”Category” column, ”G” refers to generative models, while ”D” refers to discriminative models.
  Model Category DNSMOS $\uparrow$ NISQA $\uparrow$ SpeechBERT $\uparrow$ dWER $\downarrow$ WavLM Sim $\uparrow$ Wespeaker Sim $\uparrow$
SIG BAK OVL
Mixture - 3.383 3.098 2.653 2.453 0.572 0.792 0.847 0.759
Spex+[3] D 3.378 3.771 3.000 3.029 0.840 0.213 0.964 0.922
WeSep[24] D 3.563 3.931 3.228 4.041 0.922 - 0.991 -
TSELM-L [16] G 3.550 4.084 3.228 4.029 0.808 0.275 0.908 0.627
AnyEnhance [12] G 3.638 4.066 3.353 4.277 0.735 - 0.914 -
LauraTSE G 3.609 4.084 3.336 4.333 0.908 0.159 0.974 0.876
Our AR decoder-only LM aims to predict the joint probability distribution of the target speech $\hat{s}$ conditioned on reference speech and the mixture speech embeddings according to the probability chain rule:
$P_{\theta}(\hat{s}\mid r,m)=\prod_{i\leq T}{P_{\theta}(\hat{s_{i}}\mid\hat{s}_{1:i-1},r,m)}$
where $T$ denotes the length of the output signal, and $\theta$ denotes our model parameters.
The input sequence to the AR decoder-only LM is formatted as $[<\text{bos}>,r,<\text{sep}>,m,<\text{tse}>]$, where $bos$ is a learnable token representing the start of the sentence, and $sep$ is a token that separates the reference embeddings and the mixture embeddings. $tse$ is used to split the given input and the generated output. The model is trained to generate the discrete codec embeddings of the target speech, denoted as $s$, using causal attention over $s$ to ensure the auto-regressive property. During inference, the model generates $s$ token by token, conditioned on the past tokens and the input embeddings $r$ and $m$.
The training objective is the cross-entropy loss between the predicted tokens and the ground-truth discrete tokens obtained from a neural audio codec. Specifically, we use the first $n_{q}$ layers of the codec’s Residual Vector Quantization (RVQ) as the target output. Once these $n_{q}$ layers of discrete tokens are predicted, they are mapped to their corresponding embeddings via the neural audio codec’s own embedding table. These $n_{q}$ discrete embeddings are then summed to one single embedding as the output. as well as the condition for the next token prediction.
By choosing $n_{q}$ to be smaller than the total number of RVQ layers, we simplify the modeling task, allowing the AR decoder to focus on generating coarse-grained representations of the target speech while still preserving intelligibility.
II-C Vocoder
[…]
III Experiments setup
III-A Dataset
We train our models on the 460-hour clean speech subset of LibriSpeech [25]. The training data is generated on-the-fly by mixing speech samples with a relative SNR randomly sampled between 0 and 5 dB. For cross-validation, we use the clean dev set from Libri2Mix [26]. During both training and evaluation, the reference audio is clipped to the first 5 seconds. For evaluation, we use the clean test set of Libri2Mix. Reference utterances are randomly selected for each target speaker to simulate realistic target speaker extraction conditions.
III-B Model Details
We adopt LauraGPT [18] as the backbone for our AR decoder-only LM, and use FunCodec [22] as the neural audio codec. For encoding, we apply a hop size of 256 and a window size of 512 to both the reference and mixture speech. The conformer encoder consists of 6 layers, each with 8 attention heads and a hidden dimension of 512. The decoder-only transformer has 10 layers, 8 attention heads, and a hidden size of 512. The encoder-only transformer used in the vocoder also comprises 6 layers with 8 attention heads and a 512-dimensional feature space.
Our LauraTSE model is trained from scratch. It contains a total of 77M parameters, with 36M allocated to the decoder-only transformer. We train the model using the Adam optimizer with an initial learning rate of $1\times 10^{-3}$. A warm-up scheduler with 10,000 warm-up steps is employed, and the learning rate is halved if the evaluation performance does not improve within 3 consecutive epochs. Training is conducted on 16 GPUs, each equipped with 32GB of memory, for a total of 100 epochs.
III-C Evaluation Metrics
Traditional metrics such as PESQ [27], SI-SNR, and STOI [28] are not used due to the potential misalignment between the generated waveform from vocoders and the original waveform [19]. Therefore, we use:
  • DNSMOS [29]: a reference-free metrics that has three scores from 1 to 5: SIG, BAK, and OVRL, representing the signal quality, background noise and the overall quality, respectively.
  • NISQA [30]: Another reference-free metric that predicts an overall quality score between 1 and 5 for the generated speech.
  • SpeechBERTScore [31]: A semantic similarity metric based on BERTScore computed over self-supervised speech representations. We use the HuBERT-base model to extract the features for comparison between the generated and reference speech.
  • Differential Word Error Rate (dWER) [32]: This metric computes the word error rate between the generated and reference speech using an ASR model. We employ the base model of Whisper [33] for this evaluation.
  • Speaker Similarity: This metric measures the speaker similarity between the output and ground-truth speech via cosine similarity over high-dimensional embeddings. We use two models for this task: WavLM [^2] https://huggingface.co/microsoft/wavlm-base-plus-sv and the Resnet_221LM model from WeSpeaker [34].
TABLE II: Ablation studies of LauraTSE.
  Model DNSMOS $\uparrow$ NISQA $\uparrow$ SpeechBERT $\uparrow$ dWER $\downarrow$ WavLM Sim $\uparrow$ Wespeaker Sim $\uparrow$
SIG BAK OVL
Base (Nq-2) 3.626 4.102 3.360 4.241 0.880 0.241 0.965 0.847
Nq-1 3.604 4.100 3.339 4.201 0.861 0.266 0.958 0.830
Nq-3 3.618 4.095 3.350 4.270 0.880 0.235 0.967 0.853
Ref output 3.588 4.071 3.318 4.182 0.859 0.237 0.962 0.851
Discrete IO 3.562 4.035 3.268 3.940 0.810 0.421 0.952 0.835
WavLM input 3.507 3.951 3.137 3.220 0.792 0.447 0.86 0.633
Nemo Conformer 3.621 4.101 3.350 4.243 0.734 1.567 0.929 0.765
III-D Baseline models
We compare LauraTSE with several recent baselines. First, we include Spex+ [3], a discriminative TSE model trained on Libri2Mix. We also compare it with the BSRNN model from WeSep [24] using the results provided in[12]. Additionally, we include AnyEnhance [12], a multi-task model based on masked generative modeling, and we compare with the TSE results.
IV Results and Discussions
Table I presents the overall results on the Libri2Mix Clean test set. LauraTSE achieves comparable speech quality to AnyEnhance, while outperforming it in both speaker similarity and semantic similarity. Notably, LauraTSE is trained on only 460 hours of data, whereas AnyEnhance is trained on 5000 hours, raising the question of whether multi-task learning is always superior to task-specific models for a particular objective. Compared with TSELM [16], which uses discrete tokens from WavLM, our model leverages neural audio codec representations that better preserve speaker identity, addressing the issue of low speaker similarity. Additionally, LauraTSE outperforms discriminative baselines such as Spex+ [3] and WeSep [24] in terms of speech quality, while maintaining competitive performance in semantic similarity and speaker consistency. Table II shows the results of ablation studies using models trained on Libri2Mix. ”Base” refers to the proposed lauraTSE model. ”Nq-” refers to the output number of layers of the AR model. Changing $n_{q}$ from 1 to 3 results in minimal performance differences, suggesting that even a small number of coarse layers may provide sufficient information for the AR model.
Auto-regressive decoder-only model like LauraGPT [18] conducts tasks like SE where the output is strictly aligned with the input length. Therefore, following their approach, we have formatted the decoder-only input to be $[<\text{bos}>,r,m,<\text{tse}>]$. Unlike the original method, which only produces the clean speech, this approach concatenates the reference and the mixture speech into a single input sequence, expecting the model to generate an output containing both the reference speech and the enhanced speech. This method is referred to as the ”Ref output.” During inference, we retain only the portion of the output corresponding to the mixture, constrained by the length of the reference speech. This approach yields results similar to the original one but has led to some undesirable cases where the output length is zero. In contrast, our original method demonstrates that the output sequence can be just the clean speech and does not need to align with the continuous input condition sequence.
Following the approach in SpeechX [17], we conduct experiments where the input sequences consist of discrete token embeddings rather than continuous log-mel spectrograms, referred to as ”Discrete IO.” For both the input reference speech and the mixture speech, instead of using continuous log-mel spectrograms, we utilize the first two layers of audio codec. We then use two learnable embedding matrices to embed the discrete tokens and summarize these embeddings into a single embedding. This single embedding is then fed into the AR model. We observe that these discretized representations perform worse than the continuous approach. One possible reason for this could be that the neural audio codec representations may not be optimal for usage with our small-scale AR model. Additionally, since the audio codec is typically trained on clean speech, it might miss crucial information from the mixture. A potential direction for future work could involve developing audio codecs that can effectively handle mixture speech.
We also utilize WavLM features instead of the neural audio codec as the embedding features. We use the output from the 6th hidden layer of WavLM as input features for both the reference and the mixture. Additionally, we apply the concatenation technique used in TSELM [16] for mixture representations. Following the approach in SELM [19], the output of the AR model is the Kmeans discrete representations of the target speech. After obtaining the target discrete embeddings, we use a conformer detokenizer to reconstruct the continuous embeddings, as done in SELM [19]. Similar as [16], this approach results in poor speaker similarity, likely due to the discretization process that loses speaker information. In addition, we observe that the semantic similarity is also low, raising the question of how to effectively integrate self-supervised models within AR models.
It has been shown that using a pretrained ASR conformer is beneficial for the Speaker Verification task [35]. We apply the pretrained Nemo Conformer from the Nemo Toolkit [36] to the input mixture and reference (denoted as ”Nemo Conformer”). However, this approach does not outperform the original one and yields poor dWER results. Some problematic cases lead to infinite looping. One possible reason could be the batch normalization problem as stated in [18], while another might be the limited amount of training data. Further research is needed to address these issues.
V Conclusion
We propose LauraTSE, an Auto-Regressive (AR) Decoder-Only language model (LM) designed for Target Speaker Extraction. It consists of a small-scale AR decoder-only LM that predicts the coarse-grained information of the target speech using the continuous representations of the reference and the mixed speech, and a one-step encoder-only LM that captures the fine-grained acoustic details. Extensive experiments demonstrate the model’s capability in promising speech quality, intelligibility, and speaker similarity.
Acknowledgment
This research is funded by DKU foundation project ”Emerging AI Technologies for Natural Language Processing”. Many thanks for the computational resource provided by the Advanced Computing East China Sub-Center.